Three samples of gas each exert 740. mm Hg in separate 1 L containers. What pressure do they exert if they are all placed in a single 2 L container ? Show complete solution


1
Expert's answer
2021-11-08T08:12:19-0500

P1V1= P2V2

740×1= 2×V2

740= 2V2

V2= 740/2

= 370mmHg